# Robustness of Decentralised Learning to Nodes and Data Disruption

### v1

@@ 5.4 Knowledge persistence
First, we analyse the overall performance of the system, considering all nodes altogether. To this end, in Figure 7(a) we show the average accuracy of the entire system, as defined in Equation 2. As we can see from Figure 7(a), there is a noticeable gap in accuracy with respect to the baseline. Furthermore, from the zoom inset, we can observe that, despite the curves being closely grouped together, there is a gap between the system’s performances for the different accuracy thresholds. This can be better analysed by looking at the percentage difference (Figure 7(b)). With increasing accuracy threshold, the distance from the baseline decreases, and overall, it ranges anyway around 4 and 8%. Also, in this case, we notice the slightly lower performance of Case 2 with respect to Case 1.
@@
Next, we focus individually on the two boundary cases of completely isolated nodes, and on the case of the biggest cluster. In 8(a), we show the mean accuracy of the isolated nodes for different accuracy thresholds. We can observe several interesting features. First, we confirm the drop in accuracy after disruption suffered at any accuracy threshold (higher than 0). This means that, for isolated nodes, it is fundamental to get access to knowledge extracted from data residing on other nodes and that they cannot fully compensate for the effect of a disruption using local data only. On the other hand, it is also interesting to observe that isolated nodes, at any accuracy threshold, achieve higher accuracy with respect to the case where they started in isolation (accuracy threshold equal to 0). This means that, despite the observed drop, some knowledge acquired thanks to collaborative learning still persists after disruption and “stays there” for a long time. In other words, this knowledge persists even in the absence of a large representation of labels in data residing locally. The difference metric (8(b) shows that the drop in accuracy ranges between 10 and 20% depending on the accuracy threshold. All in all, Figure 8 shows a significant persistence of knowledge after disruption even at isolated nodes, thanks to the fact that they had the chance of being exposed to the collaborative learning process before disruption, even though some loss of knowledge is unavoidable. The same behaviour can also be observed for small clusters and vanishes only when clusters contain an amount of data sufficient to recover from the loss of global knowledge diffusion implied from the time of disruption.
@@
Figure 9: Knowledge persistence analysis (largest cluster). (a) Mean accuracy over all nodes in the largest cluster. (b) Mean accuracy difference with respect to baseline over all nodes in the largest cluster.
@@
Figure 9 illustrates the average accuracy attained by nodes within the largest cluster and the percentage difference compared to the baseline. Intuitively, these nodes are less affected by disruptions, which is supported by the results. Specifically, we observe the following main findings. Firstly, large clusters demonstrate resilience to disruptions regardless of the accuracy thresholds, as evidenced by the nearly overlapping accuracy curves. This indicates that the timing of the disruption minimally impacts the performance of clusters. Secondly, there is a slight difference in performance compared to the baseline (approximately 1%), suggesting that large clusters can effectively mitigate the effects of disruptions quite efficiently. One particular aspect to note in this case is that the accuracy in Case 1 is maximised when the disruption occurs at time 0, and this effect is even more evident for earlier communication rounds. In other words, in Case 1, nodes of the largest cluster do not benefit at all from the presence of the other nodes in the network, which are actually detrimental to them. Albeit counter-intuitive, this behaviour can be explained also based on the analysis we have carried out in [14]. Remember that in Case 1 the most central nodes (which undergo disruption) do not have data of their own, but they just average and forward models received from neighbours without any local retraining. Particularly early on in time, these nodes receive very heterogeneous models (as they are central, they typically have many neighbours, see Section 4.2) and therefore, the average model they compute is not very accurate unless it has a chance of being retrained on local data. As a side effect, when disruption occurs after time 0, central nodes inject in the largest (after disruption) cluster “noisy” models, which do not contribute to, yet work against, the overall accuracy achieved by nodes in this cluster.
@@
Considering the results presented in this section altogether, we can observe a significant persistence of knowledge after disruption in all cases. Even in the most unfavourable conditions (i.e., isolated nodes) the accuracy achieved sufficiently long after the disruption is significantly higher than what would be achieved if those nodes were never exposed to the decentralised learning process (i.e., the case “accuracy threshold 0”). There is clearly a loss of accuracy with respect to the case where no disruption occurs, but this is rather limited (in the order of 20% maximum in our experiments). Moreover, if nodes remain grouped in sufficiently large clusters, the decentralised learning process is extremely robust, and, if sufficient time is allowed after disruption, the achieved accuracy is basically indistinguishable from the case when no disruption occurred.
@@ 5.5 The effect of non-IID data
In the scenarios considered so far, nodes have (as summarized in Table 1) either the same amount of images per class or none at all (disrupted nodes in Case 1), so the data distribution was IID assuming that data were present. This implied that each node with data would contribute equally to the learning process. In this final set of results, we want to focus, instead, on the effects of an unequal data distribution. As we explained in Section 4.4, in Case 3, data labels are divided into two classes, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Images belonging to $\mathcal{L}_{1}$ are distributed in an IID fashion across all nodes by simply splitting the approximately 6,000 images among all nodes (hence, each node gets around 60 images per class). Instead, only 10, 20 or 30 images belonging to $\mathcal{L}_{2}$ are assigned to the nodes surviving the disruption, while all the others are split equally among disrupted nodes, which end up with approximately 500 images per $\mathcal{L}_{2}$-class. Thus, in Case 3, disrupted nodes are disproportionately better at classifying $\mathcal{L}_{2}$-images hence losing them after the disruption is expected to significantly impact the learning process. Before we continue, note that in Case 3 we had to consider higher accuracy thresholds than in the previous cases (87%, 90%, 92%). This is because in this Case 3 configuration, there are much more data around in the network overall, so accuracies of 80% are basically reached in the very first communication rounds.
Figure 10 displays the accuracy over time for Case 3. Let us begin with the most challenging configuration (top row), where the surviving nodes after the disruption only possess 10 images for each $\mathcal{L}_{2}$-label (those ranging from 5 to 9 in the MNIST dataset). When the disruption occurs at time zero, the spread of curves is quite broad, indicating that some nodes lack the ability to learn solely based on local data. As the disruption is delayed, the spread narrows, indicating that even the most isolated nodes post-disruption can benefit from collaborative training that occurred before disruption, despite having very few local $\mathcal{L}_{2}$-images. Note, in the inset of the figure, that more peripheral nodes experience a drop in accuracy after the disruption, as observed in previous cases. However, the accuracy remains significantly higher than when there is no collaboration at all (corresponding to an accuracy threshold of 0%). With an increase in the number of local images (second and third rows), we observe an overall facilitation in the learning process, as expected. In this scenario, the timing of the disruption plays a less significant role, as the local data suffice to compensate for the loss of collaborators.
Zooming in on the differences between isolated nodes and the largest connected component after the disruption (Figure 11), we confirm that nodes that become completely disconnected post-disruption can capitalise on the knowledge received and assimilated before the disruption. Collaborative learning benefits them most when they have very few local training images initially. Large clusters (bottom row of Figure 11) suffer much less from the disruption because they can effectively pool their knowledge thanks to the remaining graph connectivity. However, similar to previous cases, later disruptions and an increase in local images also alleviate the burden on their learning process. These findings are clearly confirmed when looking at the differences in accuracy (Figure 12).
The results for Case 3 confirm the robustness of the decentralised learning process to network disruptions: as long as some knowledge is available and accessible through the communication graph, compensating for the loss of collaborators – even excellent ones like in Case 3 – is feasible and the loss in accuracy with respect to the baseline of no disruption is drastically mitigated. More than other, nodes that lose all collaborators benefit significantly from the accumulation of deep knowledge circulated before the disruption.
@@
Figure 10: Accuracy curves for each active node in Case 3. Top to bottom: increasing value of the number of images in the $\mathcal{L}_{2}$-label. From left to right: increasing accuracy threshold condition for the disruption. The curves are coloured based on the size of the cluster to which the corresponding node belongs. The inset zoom displays the clusterization of accuracy levels.
@@
Figure 11: Mean accuracy over isolated nodes (top) and over nodes belonging to the largest cluster (bottom), for different disruption thresholds.
@@
Figure 12: Accuracy difference with respect to baseline for: all surviving nodes (top row), isolated nodes (middle row), large cluster (bottom row), for different disruption thresholds.
6 Conclusions
@@
The most important finding we can highlight is that decentralised learning appears to be remarkably robust to all types of disruptions considered, as long as surviving nodes hold a sufficient fraction of representative data to sustain the learning process after disruption. For nodes belonging to large clusters (after disruption) the loss of accuracy is negligible compared to the case when no disruption occurred. For isolated nodes, the loss of accuracy is larger, but we have never observed it exceeding 20% with respect to the case of no disruption. In all cases, we have shown that if surviving nodes are part of the decentralised learning process for some time before disruption occurs, then they can retain most of the knowledge acquired before disruption and achieve a much higher accuracy than in the case when the network starts already in the same configuration as after disruption.
@@ Acknowledgments
This work was partially supported by the H2020 HumaneAI Net (952026) and by the CHIST-ERA-19-XAI010 SAI projects. C. Boldrini’s and M. Conti’s work was partly funded by the PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - “FAIR”, A. Passarella’s work was partly funded by the PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000001 - “RESTART”, both funded by the European Commission under the NextGeneration EU programme.

### v2

@@ 5.4 Knowledge persistence
First, we analyse the overall performance of the system, considering all nodes altogether (Figure 8). To this end, in Figure 8(a) we show the average accuracy of the entire system, as defined in Equation 2. As we can see from Figure 8(a), there is a noticeable gap in accuracy with respect to the baseline. Furthermore, from the zoom inset, we can observe that, despite the curves being closely grouped together, there is a gap between the system’s performances for the different accuracy thresholds. This can be better analysed by looking at the percentage difference (Figure 8(b)). With increasing accuracy threshold, the distance from the baseline decreases, and overall, it ranges anyway around 4 and 8%. Also, in this case, we notice the slightly lower performance of Case 2 with respect to Case 1. Next, we focus individually on the two boundary cases of completely isolated nodes (Figure 9), and on the case of the biggest connected component (Figure 10).
@@
We start with isolated nodes. In 9(a), we show the mean accuracy of the isolated nodes for different accuracy thresholds. We can observe several interesting features. First, we confirm the drop in accuracy after disruption suffered at any accuracy threshold (higher than 0). In A, we discuss additional experiments that show that this drop is due to overfitting on the local data after disruption. Regardless of overfitting, this means that, for isolated nodes, it is fundamental to get access to knowledge extracted from data residing on other nodes and that they cannot fully compensate for the effect of a disruption using local data only. On the other hand, it is also interesting to observe that isolated nodes, at any accuracy threshold, achieve higher accuracy with respect to the case where they started in isolation (accuracy threshold equal to 0). This means that, despite the observed drop, some knowledge acquired thanks to collaborative learning still survives after disruption and persists for a long time. In other words, this knowledge persists even in the absence of a large representation of labels in data residing locally. The difference metric (9(b) shows that the drop in accuracy ranges between 10 and 20% depending on the accuracy threshold. All in all, Figure 9 shows a significant persistence of knowledge after disruption even at isolated nodes, thanks to the fact that they had the chance of being exposed to the collaborative learning process before disruption, even though some loss of knowledge is unavoidable. The same behaviour can also be observed for small connected components and vanishes only when connected components contain an amount of data sufficient to recover from the loss of global knowledge diffusion implied from the time of disruption.
@@
Figure 10: Knowledge persistence analysis (largest connected component). (a) Mean accuracy over all nodes in the largest connected component. (b) Mean accuracy difference with respect to baseline over all nodes in the largest connected component.
@@
Figure 10 illustrates the average accuracy attained by nodes within the largest connected component and the percentage difference compared to the baseline. Intuitively, these nodes are less affected by disruptions, which is supported by the results. Specifically, we observe the following main findings. Firstly, large connected components demonstrate resilience to disruptions regardless of the accuracy thresholds, as evidenced by the nearly overlapping accuracy curves. This indicates that the timing of the disruption minimally impacts the performance of connected components. Secondly, there is a slight difference in performance compared to the baseline (approximately 1%), suggesting that large connected components can effectively mitigate the effects of disruptions quite efficiently. One particular aspect to note in this case is that the accuracy in Case 1 is maximised when the disruption occurs at time 0, and this effect is even more evident for earlier communication rounds. In other words, in Case 1, nodes of the largest connected component do not benefit at all from the presence of the other nodes in the network, which are actually detrimental to them. Albeit counter-intuitive, this behaviour can be explained also based on the analysis we have carried out in [8]. Remember that in Case 1 the most central nodes (which undergo disruption) do not have data of their own, but they just average and forward models received from neighbours without any local retraining. Particularly early on in time, these nodes receive very heterogeneous models (as they are central, they typically have many neighbours, see Section 4.2.1) and therefore, the average model they compute is not very accurate unless it has a chance of being retrained on local data. As a side effect, when disruption occurs after time 0, central nodes inject in the largest (after disruption) connected component “noisy” models, which do not contribute to, yet work against, the overall accuracy achieved by nodes in this connected component.
@@
Take-home: Considering the results presented in this section altogether, we can observe a significant persistence of knowledge after disruption in all cases. Even in the most unfavourable conditions (i.e., isolated nodes) the accuracy achieved sufficiently long after the disruption is significantly higher than what would be achieved if those nodes were never exposed to the decentralised learning process (i.e., the case “accuracy threshold 0”). There is clearly a loss of accuracy with respect to the case where no disruption occurs, but this is rather limited (in the order of 20% maximum in our experiments). Moreover, if nodes remain grouped in sufficiently large connected components, the decentralised learning process is extremely robust, and, if sufficient time is allowed after disruption, the achieved accuracy is basically indistinguishable from the case when no disruption occurred.
@@ 5.5 The effect of non-IID data
In the scenarios considered so far, nodes have (as summarized in Table 1) either the same amount of images per class or none at all (disrupted nodes in Case 1), so the data distribution was IID assuming that data were present. This implied that each node with data would contribute equally to the learning process. In this final set of results, we want to focus, instead, on the effects of an unequal data distribution. As we explained in Section 4.3, in Case 3, data labels are divided into two classes, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, and we considered two different sub-scenarios regarding their distribution. In both sub-scenarios, the nodes will be divided into two groups, G1 and G2, based on the amount of images belonging to $\mathcal{L}_{2}$ they are assigned. $\mathcal{L}_{1}$ images are distributed in an IID fashion across all nodes by simply splitting the approximately 6,000 images among all nodes (hence, each node gets around 60 images per class). Instead, $\mathcal{L}_{2}$ images are distributed as follows: only 10, 20 or 30 images belonging to $\mathcal{L}_{2}$ are assigned to the G1 nodes, while all the others are split equally among G2 nodes. In the first sub-scenario, G2 nodes coincide with the disrupted nodes, which end up with approximately 500 images per $\mathcal{L}_{2}$-class. Thus, in Case 3.1, disrupted nodes are disproportionately better at classifying $\mathcal{L}_{2}$-images and losing them after the disruption is expected to significantly impact the learning process. In the second sub-scenario, G2 nodes will be the 10 nodes with the lowest structural hole score, meaning that, after the disruption, among the surviving nodes there will be ones that preserve more knowledge on the $\mathcal{L}_{2}$-classes. Before we continue, note that in Case 3 we had to consider higher accuracy thresholds than in the previous cases (87%, 90%, 92%). This is because in this Case 3 configuration, there are much more data around in the network overall, so accuracies of 80% are basically reached in the very first communication rounds.
5.5.1 Case 3.1: Disrupted nodes with more knowledge
@@
Figure 11: Accuracy curves for each active node in Case 3.1. Top to bottom:increasing accuracy threshold condition for the disruption. From left to right: increasing value of the number of images in the $\mathcal{L}_{2}$-label. The curves are coloured based on the size of the connected component to which the corresponding node belongs. The inset zoom displays the clusterization of accuracy levels.
Figure 11 displays the accuracy over time for Case 3.1. Let us begin with the most challenging configuration (first column), where the surviving nodes after the disruption only possess 10 images for each $\mathcal{L}_{2}$-label (those ranging from 5 to 9 in the MNIST dataset). When the disruption occurs at time zero, the spread of curves is quite broad, indicating that some nodes lack the ability to learn solely based on local data. As the disruption is delayed, the spread narrows, indicating that even the most isolated nodes post-disruption can benefit from collaborative training that occurred before disruption, despite having very few local $\mathcal{L}_{2}$-images. Note, in the inset of the figure, that more peripheral nodes experience a drop in accuracy after the disruption, as observed in previous cases. However, the accuracy remains significantly higher than when there is no collaboration at all (corresponding to an accuracy threshold of 0%). With an increase in the number of local images (second and third columns), we observe an overall facilitation in the learning process, as expected. In this scenario, the timing of the disruption plays a less significant role, as the local data suffice to compensate for the loss of collaborators.
Zooming in on the differences between isolated nodes and the largest connected component after the disruption (Figure 12), we confirm that nodes that become completely disconnected post-disruption can capitalise on the knowledge received and assimilated before the disruption. Collaborative learning benefits them most when they have very few local training images initially. The largest connected component (second column of Figure 12) suffers much less from the disruption because it can effectively pool its nodes’ knowledge thanks to the remaining graph connectivity. However, similar to previous cases, later disruptions and an increase in local images also alleviate the burden on their learning process. These findings are clearly confirmed when looking at the differences in accuracy (Figure 13).
@@
Figure 12: Mean accuracy in Case 3.1 over isolated nodes (left) and over nodes belonging to the largest connected component (right), for different disruption thresholds and number of $\mathcal{L}_{2}$-images.
@@
Figure 13: Accuracy difference in Case 3.1 with respect to baseline for: all surviving nodes (top row), largest connected component (middle row), isolated nodes (bottom row), for different disruption thresholds.
Take-home: The results for Case 3.1 confirm the robustness of the decentralised learning process to network disruptions: as long as some knowledge is available and accessible through the communication graph, compensating for the loss of collaborators – even excellent ones like in Case 3.1 – is feasible and the loss in accuracy with respect to the baseline of no disruption is drastically mitigated. More than other, nodes that lose all collaborators benefit significantly from the accumulation of deep knowledge circulated before the disruption.
5.5.2 Case 3.2: Peripheral nodes with more knowledge
Figure 14 shows the accuracy over time for Case 3.2, which follows the same behavior as that of Case 3.1, with the narrowing of the curves as the disruption is delayed proving the impact of collaborative learning. A notable distinction arises in Case 3.2, where some of the blue curves — which represent isolated nodes — exhibit higher accuracy levels compared to others. This phenomenon can be attributed to the fact that some of the isolated nodes belong to the G2 group (with a high number of samples for $\mathcal{L}_{2}$-data). Figure 15 shows the accuracy progression over time for isolated nodes when G1 nodes have access to 10 images of the $\mathcal{L}_{2}$ class, with G2 nodes distinctly highlighted. Moreover, going from left to right, we can observe how the distance in performance of the G1 and G2 isolated nodes evolves w.r.t the delay in the disruption. We can see that the G1 nodes eventually reach comparable accuracy level of the G2 nodes. This convergence in performance clearly indicates that delaying the disruption of collaborative learning benefits the G1 isolated nodes, allowing these nodes to compensate for their smaller amount of $\mathcal{L}_{2}$ data. This is the effect of knowledge persistence. The extended interaction period helps G1 nodes to leverage the collaborative learning, improving their ability to perform well even after their isolation.
Figure 14: Accuracy curves for each active node in Case 3.2. Top to bottom: increasing accuracy threshold condition for the disruption. From left to right: increasing value of the number of images in the $\mathcal{L}_{2}$-label. The curves are coloured based on the size of the connected component to which the corresponding node belongs. The inset zoom displays the clusterization of accuracy levels.
Figure 15: Accuracy over time in Case 3.2 for the isolated nodes. Coloring is based on their group membership. From left to right: increasing accuracy threshold. Top to bottom: increasing value of the number of images in the $\mathcal{L}_{2}$-label.
5.5.3 Comparison between Case 3.1 and Case 3.2
Figure 16: Accuracy distance of nodes in the largest (after disruption) connected component between Case 3.1 and Case 3.2. A positive difference implies that the accuracy of Case 3.2 is higher than that of Case 3.1, and vice versa. The dotted line represents the equivalence between the two cases.
The major difference between the two sub-scenarios is that some G2 nodes will survive the disruption in Case 3.2. This difference influences the behaviour of the system, as seen when comparing the accuracy levels of nodes within the largest connected component (LCC) between Case 3.1 and Case 3.2, as shown in Figure 16. To this purpose, we use Definition 4.3. Initially, while the network remains fully connected, Case 3.1 exhibits higher performance than Case 3.2, reflected in a negative distance between the accuracies. This suggests that during the collaborative phase, nodes in Case 3.1 are able to benefit more from the shared learning process, because more knowledge is placed on the nodes that are more central, hence the spreading of knowledge is faster. However, as time progresses and the network experiences disruption, the trend shifts. In the long term, nodes in Case 3.2 achieve higher accuracy levels, mainly due to the sustained presence of G2 nodes, which ensures a greater overall availability of data within the network. Moreover, as the number of $\mathcal{L}_{2}$ images accessible locally to G1 nodes increases (from left to right), the gap in accuracy between the two cases narrows (i.e., the difference approaches zero). This indicates that the collaborative learning process within the LCC helps to mitigate the initial disadvantage faced by G1 nodes due to their smaller data sets. Essentially, the prolonged interaction with other nodes allows G1 nodes to better utilize the data available through collaboration, partially compensating for the lack of direct access to $\mathcal{L}_{2}$-data. This is especially notable by looking at the blue curve, where the system starts as already disrupted in both cases.
Take-home: In the long run, after a disruption, the learning process is more effective when more data are retained within the surviving network, even if they reside on non-central nodes. Consequently, in Case 3.2, the final accuracy is higher than in Case 3.1. However, the difference is not large and decreases the later the disruption occurs and the larger the local datasets become. Additionally, the same amount of data stored in central nodes rather than in peripheral nodes accelerates the learning process, as demonstrated by the superior initial (pre-disruption) performance of Case 3.1 compared to Case 3.2.
6 Conclusions and future work
@@
The most important finding we can highlight is that decentralised learning appears to be remarkably robust to all types of disruptions considered, as long as surviving nodes hold a sufficient fraction of representative data to sustain the learning process after disruption. For nodes belonging to large connected components (after disruption) the loss of accuracy is negligible compared to the case when no disruption occurred. For isolated nodes, the loss of accuracy is larger, but we have never observed it exceeding 20% with respect to the case of no disruption. In all cases, we have shown that if surviving nodes are part of the decentralised learning process for some time before disruption occurs, then they can retain most of the knowledge acquired before disruption and achieve a much higher accuracy than in the case when the network starts already in the same configuration as after disruption.
@@
This work, along with [22], marks an initial step toward a deeper understanding of the interplay between the topology and dynamics of complex systems and the learning processes they support. However, many important research questions remain unexplored. A natural extension of this work would be to explore more sophisticated learning strategies beyond DecAvg, and to apply the framework to more challenging datasets than the baseline MNIST. Additionally, the network topologies studied could be expanded to include more real-world graphs, as well as synthetic ones that are less representative of real systems (e.g., regular graphs). Analyzing these diverse topologies could help establish upper and lower bounds on robustness as function of different network classes. From the disruption standpoint, this work focused on the failure of entire nodes at a specific moment in time. However, future research could explore more dynamic failure models in which failures occur gradually over time, potentially affecting edges rather than nodes, or causing nodes to fail sequentially instead of abruptly all at once. We also assume that nodes do not reconnect once disrupted, whereas in real-world scenarios, disruptions might be temporary. A deeper analysis could investigate the impact of temporary disruptions, especially in heterogeneous computing environments where nodes operate at varying speeds, since the disruption of faster versus slower nodes could affect the learning process differently. Another important direction is the theoretical derivation of robustness properties, rather than solely inferring them from experiments, which could offer a more formal understanding of the system’s behavior. Additionally, an orthogonal yet critical research avenue is robustness against malicious threats, rather than just failures, a topic that remains largely unexplored in the current literature on fully decentralised learning.
@@ Acknowledgments
This work was partially supported by the H2020 HumaneAI Net (952026) and by the CHIST-ERA-19-XAI010 SAI projects. C. Boldrini’s and M. Conti’s work was partly funded by the PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - “FAIR”, A. Passarella’s and L. Valerio’s work was partially supported by the European Union - Next Generation EU under the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3, CUP B53C22003970001, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”). J. Kertész acknowledges support also from ERC grant No. 810115-DYNASNET.
Appendix A Overfitting on isolated nodes
When looking at the accuracy curves of isolated nodes (Figure 9), we often noticed a drop in accuracy after a disruption. This drop observed in isolated nodes can be traced back to the limited number of images per class available in the individual datasets. As indicated in Table 1, the individual datasets contain either 7 images per class (Case 1) or 6 images per class (Case 2). This small number of images per class may lead to overfitting when nodes become completely isolated. With such limited data, after disruption, the local model begins to memorize the specific properties of the few available images rather than learning general patterns, which gradually reduces its ability to generalize effectively to new data (i.e., those in the test set). To prove this, we conducted an additional set of experiments where we doubled the number of images per class. The results, presented in Figure 17, show the mean accuracy curves for all isolated nodes in both Case 1 and Case 2, comparing the small local training dataset used in the previous experiments with larger versions. It is evident that the significant drop in accuracy observed with the smaller datasets disappears when the local data availability is increased. While a decreasing trend in accuracy remains, the overfitting problem is notably mitigated, highlighting the importance of sufficient data for generalization in isolated nodes.
Figure 17: Mean accuracy curves of the isolated nodes for Case 1 and Case 2 and counterparts with doubled local training set size.
Appendix B Karate Club Graph
As stated in Section 5, as a real-world network example, we analyzed the Karate Club Graph [25]. The Karate Club Graph exhibits clear and well-defined community structures due to the actual affiliations and conflicts within the karate club. This makes it highly representative of real community dynamics. The network is shown in Figure 18, before and after the disruption occurs. It seems that the removal of four highly central nodes destroys the community structure. Note that the colouring of nodes is based on the size of the connected component they belong to after disruption. Following the same considerations presented in Section 4.2.1, for the Karate Club graph we decided to remove a percentage of nodes that matches the LCC size ratio of the Barabasi-Albert graph. As shown in Figure 19, this amounts to the 12% which means that we are removing 4 nodes. The results obtained confirm the overall system’s behavior obtained for the BA network. In Figure 20, we show the accuracy of all the surviving (i.e., non-switched-off) nodes in Case 1 and Case 2 when the disruption happens at different accuracy thresholds. As we observed for the BA, the later the disruption happens, i.e., going from left to right, the higher the mean accuracy of the system, resulting in a narrower curve beam. Furthermore, we observe the same clusterization into different groups of similarly performing nodes, based on their connected components’ sizes. In Figure 21(a) we show the average accuracy of the entire system (as for Definition 2). There is a noticeable decrease in accuracy compared to the baseline and the zoomed inset reveals that even though the performance curves are closely clustered, there is a discernible gap between the system’s performance at different accuracy thresholds. This discrepancy becomes more evident when analyzing the percentage differences shown in Figure 21(b). As the accuracy threshold increases, the deviation from the baseline decreases, generally ranging between 4% and 9%. Additionally, we observe that Case 2 consistently exhibits slightly lower performance compared to Case 1. Focusing on the isolated nodes, Figure 22 illustrates that the knowledge obtained from the collaborative task persists, resulting in a percentage decrease in accuracy of up to 20% in the worst case (accuracy threshold of 70%). Similarly, we observe that the performance of the nodes in the largest connected component (LCC) does not differ significantly from the baseline case (Figure 23(a) and 23(b)). This consistency is due to their continuous stream of connections, which compensates for the disruption. Overall, the robustness that decentralized learning demonstrated in the Barabási-Albert (BA) network is also evident in the Karate Club Graph. The loss of accuracy compared to the scenario with no disruption is negligible within the LCC, and even for completely isolated nodes, where it remains limited to between 10% and 20%, depending on the timing of the disruption. This indicates that decentralized learning maintains a high level of performance despite significant network impairments. In addition, it seems that the removal of four highly central nodes destroys the initial community structure of the Karate Club Graph (see Figure 18), so that it does not have an impact on the learning process.
Figure 18: Karate club network before (left) and after (right) the disruption. The colouring of nodes is based on the size of the connected component they belong to after disruption, using as basis the connected components’ coloring used for the Barabasi-Albert. The black nodes in the left image are the nodes that will be switched off when the disruption condition is fulfilled.
Figure 19: Karate Club Graph’s size of the largest connected component as nodes are removed progressively. The dotted line represent the percentage chosen to be removed to account for the same LCC size of the BA network.
{adjustwidth}
-1cm-1cm
Figure 20: Accuracy curves for each surviving node in Case 1 (a) and Case 2 (b). From left to right: increasing accuracy threshold condition for the disruption. The curves are coloured based on the size of the connected component to which the corresponding node belongs after the disruption. Accuracy threshold 75% is omitted for ease of visualisation, as it only showed an intermediate behaviour between threshold 70% and 80%.
Figure 23: Knowledge persistence analysis (largest connected component). (a) Mean accuracy over all nodes in the largest connected component. (b) Mean accuracy difference with respect to baseline over all nodes in the largest connected component.